# Feedback-Enhanced Online Multiple Testing with Applications to Conformal Selection

### v1

@@ 2.2 Boosting GAI via feedback: GAIF
In our problem setup, at each time step $t$, feedback reveals the true values of $\{\theta_{j}\}_{j=1}^{t-1}$. The key idea lies in leveraging this feedback to reduce statistical slack—enhancing both the accuracy of FDP estimation and the efficiency of alpha-wealth allocation. We formalize this in the following GAI with Feedback (GAIF) framework.
@@
The GAIF procedure refers to any rule for assigning test levels $\alpha_{t}$ such that
$\widehat{\operatorname{FDP}}_{\rm{GAIF}}(t):=\frac{\sum_{j=1}^{t-1}(1-\theta_{j})\,\alpha_{j}\,\kappa(p_{j})+\alpha_{t}\,\kappa(p_{t})}{1\vee\sum_{j=1}^{t}\delta_{j}}\leq\alpha,$
where $\kappa\colon[0,1]\to\mathbb{R}_{\geq 0}$ denotes a predetermined weight function satisfying
${\mathbb{E}}\{\widehat{\operatorname{FDP}}_{{\rm GAIF}}(t)\}\geq{\mathbb{E}}\!\Bigl{\{}\tfrac{\sum_{j\leq t}(1-\theta_{j})\,\alpha_{j}}{1\vee\sum_{j\leq t}\delta_{j}}\Bigr{\}}.$ (3)
The corresponding testing procedures are summarized in Algorithm 1.
@@
0:  Target $\operatorname{FDR}$ level $\alpha$, pre-specified parameters for constructing test levels
@@
2:   Observe $p$-value $p_{t}$
3:   Update $\alpha_{t}=\alpha_{t}^{\text{GAIF}}$, constructed according to Definition (2.1)
4:   if $p_{t}\leq\alpha_{t}$ then $\delta_{t}=1$, else $\delta_{t}=0$
5:   Obtain the revealed feedback $\theta_{t}$
6:  end for
6:  Rejection set ${\mathcal{R}}=\{t:\delta_{t}=1\}$.
The GAIF framework improves statistical power through two mechanisms. The first one is leveraging feedback $\{\theta_{i}\}_{i<t}$ to improve the accuracy of FDP estimation. By incorporating feedback, GAIF can identify which previously tested hypotheses are null ($j\leq t-1$), enabling a precise contribution of past tests to the FDP and reducing conservativeness. Importantly, this refinement can be applied to almost all GAI procedures—except for LOND [Javanmard and Montanari, 2015]; see Appendix A.7.
Another mechanism is to efficiently allocate the alpha wealth by adopting the weight function $\kappa_{j}$. Unlike offline multiple testing—where a common threshold applies to all hypotheses—online procedures must determine each $\alpha_{t}$ individually. Therefore, power depends not only on accurate FDP estimation but also on how alpha-wealth is allocated over time. To improve allocation, GAIF introduces data-adaptive weight functions $\kappa(\cdot)$.
Formally, GAIF selects $\alpha_{t}$ to satisfy
@@
$\alpha_{t}\;\leq\;\frac{1}{\kappa(p_{t})}\Biggl{\{}\alpha\sum_{j=1}^{t}\delta_{j}\;-\;\sum_{j=1}^{t-1}(1-\theta_{j})\,\alpha_{j}\,\kappa(p_{j})\Biggr{\}}$ (4)
which, together with (3), provides the guiding principle for alpha-wealth allocation.
Choice of $\kappa(\cdot)$. The weight function $\kappa(\cdot)$ should satisfy:
  1. 1.
    Nonnegativity: $\kappa(u)\geq 0$ for all $u\in[0,1]$.
  2. 2.
    Null unbiasedness: $\mathbb{E}\{\kappa(U)\}=1$, where $U\sim\mathrm{Uniform}(0,1)$.
  3. 3.
    Concentration on large $p$-values: $\kappa$ places more weight on high $p$-values so that, when nulls prevail, “unused” alphaalpha-wealth is reclaimed and reallocated to future tests.
By these requirements, more $\alpha$-wealth is reserved for tests with small pp–values, increasing power whenever non-null alternatives are present. Indeed, from inequality (4) we see that, when the non-null proportion is large, for small $p_{t}$, the weighted rule produces an $\alpha_{t}$ that strictly exceeds its unweighted counterpart, thereby further boosting power.
@@
In some applications, the feedback is delayed: instead of seeing $\theta_{t-1}$ at time $t$, it arrives at time $t+d$, where $d\geq 0$ denotes the delay. At time $t$, we have $\{\theta_{i}\}_{i=1}^{t-d-1}$. Our GAIF framework can be extended to this case naturally, say we need
$\widehat{\operatorname{FDP}}_{\text{delay}}(t):=\frac{\sum_{j=1}^{t-d-1}(1-\theta_{j})\,\alpha_{j}\,\kappa(p_{j})+\sum_{j=t-d}^{t}\alpha_{j}\,\kappa(p_{j})}{1\vee\sum_{j=1}^{t}\delta_{j}}\leq\alpha.$
2.3 Two concrete approaches of GAIF: LF & SF
The remaining question is how to choose the weight function $\kappa(\cdot)$ in practice. We address this by presenting two concrete examples: LF and SF.
LF: an initial approach to improve FDP estimation
The simplest choice is setting all $\kappa_{j}(p_{j})=1$ by noting that $\theta_{t}\leq 1$, resulting in a feedback-enhanced LORD++ estimator, $\widehat{\operatorname{FDP}}_{\text{LF}}$ (denoted as LF), given as
$\widehat{\operatorname{FDP}}_{\text{LF}}(t):=\frac{\sum_{j=1}^{t-1}\alpha_{j}(1-\theta_{j})+\alpha_{t}}{\sum_{j\leq t}\delta_{j}\vee 1}.$
Enforcing $\widehat{\operatorname{FDP}}_{\text{LF}}(t)\leq\alpha$ is equivalent to $\sum_{j=1}^{t}\alpha_{j}\leq\sum_{j=1}^{t}\alpha\delta_{j}+\sum_{j=1}^{t-1}\alpha_{j}\theta_{j}$, and therefore we construct the corresponding test levels by:
$\alpha_{t}^{\text{LF}}=\gamma_{t}s_{0}+(\alpha-s_{0})\gamma_{t-\tau_{1}}\mathbb{I}\{\tau_{1}<t\}+\alpha\sum_{j:\tau_{j}<t,\tau_{j}\neq\tau_{1}}\gamma_{t-\tau_{j}}+\sum_{j:j<t}\gamma_{t-j}\alpha_{j}\theta_{j},$ (5)
where $\{\gamma_{t}\}_{t=1}^{\infty}$ is a given infinite non-increasing sequence of positive constants that sums to one, $\tau_{j}$ is the time of the $j$-th rejection, and $s_{0}>0$ is the pre-specified initial wealth. When $\theta_{j}=0$ for all $j$, this approach reduces exactly to the LORD++.
As the test level of LF $\alpha_{t}^{\text{LF}}$ is not smaller than that of LORD++, this LF procedure would be more powerful than LORD++ by fully incorporating available feedback information. Next, we show that by explicitly utilizing the observed $p$-value patterns in the design of $\kappa(\cdot)$, rather than simply setting it to the constant 1, more accurate FDP estimation and more efficient alpha-wealth allocation can be achieved.
SF: an approach to achieve adaptive $\alpha$-wealth allocation
To exploit patterns in the observed $p$-values, we propose the following concrete instantiation:
$\kappa(p_{j})=\frac{\mathbf{1}\{p_{j}>\lambda\}}{1-\lambda},$ (6)
where $\lambda$ is a user-chosen constant parameter in the interval $[0,1]$ for identifying large $p$-values.
Under a true null hypothesis $\mathbb{H}_{0j}$, $\mathbb{E}\{\kappa(p_{j})\}=1$, thereby satisfying $\operatorname{FDR}(t)\leq{\mathbb{E}}\{\widehat{\operatorname{FDP}}_{\text{GAIF}}(t)\}$. This choice ensures that any “unused” budget is returned for future allocation after screening $p_{j}>\lambda$ in the construction of the $\{\alpha_{t}\}$ sequence, while preserving valid FDR control. When $\theta_{j}=0$ for all $j$, this approach reduces exactly to the SAFFRON algorithm. We call this approach as SF.
The thresholds $\{\alpha_{t}\}$ for SF can be derived analogously to those in Ramdas et al. [2018], but with the tightened, feedback-aware constraint in Eq.(4). For $t=1$, $\alpha_{1}^{\text{SF}}=\min\{\gamma_{1}s_{0},\lambda\}$. For $t>1$,
$\alpha_{t}^{\text{SF}}=\min\left\{\lambda,s_{0}\gamma_{t-C_{0+}}+\{(1-\lambda)\alpha-s_{0}\}\gamma_{t-\tau_{1}-C_{1+}}+(1-\lambda)\alpha\sum_{j\geq 2}\gamma_{t-\tau_{j}-C_{j+}}+\sum_{j:j<t}\gamma_{t-j}\alpha_{j}\theta_{j}\right\},$ (7)
where we define $C_{j+}=C_{j+}(t)=\sum_{i=\tau_{j}+1}^{t-1}C_{i}$ and $C_{t}=\mathbb{I}\{p_{t}\leq\lambda\}$.
The weight function $\kappa(\cdot)$ in Eq.(6) is also used in SAFFRON [Ramdas et al., 2018], where it serves to estimate the non-null proportion and thereby improves FDP estimation. Our GAIF framework provides another intuitive perspective on the role of this function: $\kappa(\cdot)$ can also serve as a guiding mechanism for alpha-wealth allocation, leading to improved power compared to using constant weights. The indicator $\mathbb{I}\{p_{j}>\lambda\}$ ensures that alpha-wealth is spent only on the tests with $p_{j}\leq\lambda$: If $p_{j}\leq\lambda$, then $\kappa(p_{j})=0$. In this case, the FDP bound does not penalize that test, allowing the alpha-wealth to be “invested” there.
@@ 2.3 Two concrete approaches of GAIF: LF & SF
In principle, there may exist other concrete examples of $\kappa(\cdot)$ to optimize the alpha-wealth allocation. However, this online optimization remains an open problem and a direction for future work [Gang et al., 2021]. Meanwhile, the SF construction offers a theoretically grounded and practically appealing compromise, achieving robust empirical performance with minimal tuning burden.
@@
Figure 1 depicts the testing thresholds $\{\alpha_{t}\}$ over time $t$ for various procedures applied to Gaussian observations. It is clear that our methods yield larger thresholds after improving the gap via feedback, with $\alpha_{t}^{\text{SF}}>\alpha_{t}^{\text{SAFFRON}}$ and $\alpha_{t}^{\text{LF}}>\alpha_{t}^{\text{LORD++}}$ in average. This illustrates that the GAIF framework leverages alpha-wealth more effectively through feedback, thereby achieving higher power than the traditional GAI framework.
@@
Figure 1: Results for testing Gaussian observations. Line charts depict the average of $\alpha_{t}$ over time $t$ under various testing procedures, based on 500 replications with a non-null proportion of $\pi_{1}=0.5$. Details of the data generation process are provided in Subsection 4.1.
2.4 Theoretical guarantee under independence
Our first theoretical result states that both of LF and SF guarantees online mFDR control under the conditional super-uniformity of null $p$-values
where $\mathcal{F}_{t}:=\sigma(\delta_{1},\dots,\delta_{t};\theta_{1},\dots,\theta_{t})$, and achieves online FDR control under stronger independence and monotonicity assumptions. We say that the test level sequence $\{\alpha_{t}\}_{t\in\mathbb{N}}$ is a monotonic function of the past if, for all $t\in\mathbb{N}$, $\alpha_{t}$ is coordinatewise nondecreasing in the past decisions $\{\delta_{i}:i<t\}$ and feedback $\{\theta_{i}:i<t\}$ for the LF procedure, and, for the SF procedure, additionally coordinatewise nondecreasing in $\{C_{i}:i<t\}$, where $C_{i}=\mathbb{I}\{p_{i}\leq\lambda\}$.
Theorem 1 (Online mFDR and FDR Control for GAIF).
Let $\{\alpha_{t}\}_{t\in\mathbb{N}}$ be a sequence of test levels such that $\widehat{\operatorname{FDP}}_{\rm{GAIF}}(t)\leq\alpha$, and suppose the weight function $\kappa(\cdot)$ satisfies the condition in Definition 2.1. Then:
@@ 2.4 Theoretical guarantee under independence
    If the null $p$-values are conditionally super-uniform (8), the procedure guarantees $\operatorname{mFDR}(t)\leq\alpha$ for all $t\in\mathbb{N}$.
@@
    If the null $p$-values are independent of each other and of the non-nulls, and if the test level sequence $\{\alpha_{t}\}_{t\in\mathbb{N}}$ is a monotonic function of the past for all $t$, then we additionally have $\operatorname{FDR}(t)\leq\alpha\quad\text{for all }t\in\mathbb{N}.$
  • (c)
    In particular, the LF and SF procedures are concrete instances of GAIF and thus satisfy the mFDR and FDR control guarantees under the respective conditions in (a) and (b).
2.5 Addressing local dependence
This subsection focuses on extending our framework to settings with local dependence (as defined in Definition 2.2). For a strategy that addresses a more general dependence structure via e-values, the reader may refer to Appendix D.

### v2

@@ 2.2 Boosting GAI via feedback: GAIF
We now propose generalized alpha-investing with feedback (GAIF), which leverages revealed feedback information to improve FDP estimation. The key idea is that once some past hypotheses are known to be non-nulls, their contribution to the FDP upper bound can be reduced accordingly, leaving more $\alpha$-wealth available for future tests.
Formally, let
[…]
denote the index set of hypotheses whose labels are available by time $t$, and let $\bar{\mathcal{I}}_{t}=[t]\setminus\mathcal{I}_{t}$, where $t]=\{1,2,\dots,t\}$. Realistic scenarios can lead to diverse feedback structures. For example, feedback may be available only when a rejection is made or may arrive after a delay $d\geq 0$. Our framework encompasses a variety of practically relevant cases:
  • Full and instant feedback: $\mathcal{I}_{t}=\{j\in[t]:j\leq t-1\}$;
  • Bandit and instant feedback: $\mathcal{I}_{t}=\{j\in[t]:j\leq t-1,\delta_{j}=1\}$;
  • Full and delayed feedback: $\mathcal{I}_{t}=\{j\in[t]:j\leq t-d-1\}$, where $d$ is the delay;
  • Bandit and delayed feedback: $\mathcal{I}_{t}=\{j\in[t]:j\leq t-d-1,\delta_{j}=1\}$.
At time $t$, the available feedback reveals the true values of $\{\theta_{j}\}_{j\in\mathcal{I}_{t}}$. We leverage this information to refine the FDP estimator as follows.
@@
Define the feedback-enhanced FDP estimator as
$\widehat{\mathrm{FDP}}_{\mathrm{GAIF}}(t):=\frac{\sum_{j\in\mathcal{I}_{t}}(1-\theta_{j})\,\alpha_{j}\;+\;\sum_{j\in\bar{\mathcal{I}}_{t}}\alpha_{j}}{1\,\vee\,\sum_{j=1}^{t}\delta_{j}}.$
A procedure is said to be a GAI with Feedback (GAIF) procedure if it assigns test levels $\{\alpha_{t}\}$ such that, for every $t$,
$\widehat{\mathrm{FDP}}_{\mathrm{GAIF}}(t)\leq\alpha.$
The GAIF framework aligns with intuition. For $j\in\mathcal{I}_{t}$, feedback reveals whether the corresponding hypothesis is null or non-null, so their contribution to the FDP bound can be accounted for precisely through $(1-\theta_{j})\alpha_{j}$. For indices not yet revealed, i.e., $j\in\bar{\mathcal{I}}_{t}$, we keep the contribution $\alpha_{j}$. Thus, GAIF improves the FDP estimator relative to the no-feedback case.
Importantly, this refinement leads directly to a less conservative constraint on the test levels $\alpha_{t}$. Recall that LORD++ enforces $\alpha_{t}\leq\alpha(1\vee\sum_{j\leq t}\delta_{j})-\sum_{j\leq t-1}\alpha_{j}$. In contrast, a GAIF procedure selects $\alpha_{t}$ to satisfy
$\alpha_{t}\leq\alpha(1\vee\sum_{j=1}^{t}\delta_{j})-\sum_{j\in\mathcal{I}_{t}}\alpha_{j}(1-\theta_{j})-\sum_{j\in\bar{\mathcal{I}}_{t}\setminus t}\alpha_{j}.$ (4)
The presence of revealed feedback enlarges the admissible upper bound for $\alpha_{t}$. Indeed, since $1-\theta_{j}\leq 1$ for every $j\in\mathcal{I}_{t}$, (4) implies
$\sum_{j\in\mathcal{I}_{t}}\alpha_{j}(1-\theta_{j})+\sum_{j\in\bar{\mathcal{I}}_{t}\setminus t}\alpha_{j}\leq\sum_{j\in\mathcal{I}_{t}}\alpha_{j}+\sum_{j\in\bar{\mathcal{I}}_{t}\setminus t}\alpha_{j}=\sum_{j=1}^{t-1}\alpha_{j}.$
Therefore, the upper bound of $\alpha_{t}$ under GAIF is always at least as large as that under LORD++. In this sense, an appropriate procedure with feedback is expected to improve power by sharpening the FDP estimator and releasing more $\alpha$-wealth for future allocation.
We next present a concrete instance of GAIF as follows:
$\alpha_{t}^{\text{GAIF}}=\gamma_{t}s_{0}+(\alpha-s_{0})\gamma_{t-\tau_{1}}\mathbb{I}\{\tau_{1}<t\}+\alpha\sum_{j:\tau_{j}<t,\tau_{j}\neq\tau_{1}}\gamma_{t-\tau_{j}}+\sum_{j:j\in\mathcal{I}_{t}}\gamma_{j}\alpha_{j}\theta_{j}.$ (5)
Note that if all feedback is ignored so that $\mathcal{I}_{t}=\emptyset$, the $\alpha_{t}^{\rm GAIF}$ reduces exactly to that of LORD++. This further implies that the test levels $\alpha_{t}^{\rm GAIF}$ with the additional feedback term are never smaller than those of LORD++ whenever feedback is available.
Algorithm 1 summarizes the implementation of this concrete GAIF procedure under a general feedback regime. The only difference across settings lies in how the feedback-available set $\mathcal{I}_{t}$ is updated. We denote the resulting procedures under full and instant feedback, bandit and instant feedback, full and delayed feedback, and bandit and delayed feedback by LF, LF-BI, LF-FD, and LF-BD, respectively.
@@
0:  Target FDR level $\alpha$; pre-specified parameters for constructing test levels.
@@
2:   Observe the $p$-value $p_{t}$.
3:   Update $\mathcal{I}_{t}$ according to the specific feedback regime.
4:   Construct $\alpha_{t}=\alpha_{t}^{\mathrm{GAIF}}$ using Definition 2.1 with the current $\mathcal{I}_{t}$.
5:   if $p_{t}\leq\alpha_{t}$ then set $\delta_{t}=1$; else set $\delta_{t}=0$.
6:   Reveal feedback by obtaining $\theta_{j}$ for all $j\in\mathcal{I}_{t+1}\setminus\mathcal{I}_{t}$.
7:  end for
7:  Rejection set $\mathcal{R}=\{t:\delta_{t}=1\}$.
2.3 Theoretical guarantee under independence
The following result states that GAIF procedures guarantee online mFDR control under conditional super-uniformity of null $p$-values
where the enlarged filtration $\mathcal{F}_{t-1}:=\sigma(\delta_{1:t-1};\{\theta_{j}\}_{j\in{\mathcal{I}}_{t}})$. Under stronger independence and monotonicity assumptions, GAIF further achieves online FDR control. Here, the test level sequence $\{\alpha_{t}\}_{t\in\mathbb{N}}$ is said to be a monotonic function of the past if, for all $t\in\mathbb{N}$, $\alpha_{t}$ is coordinate-wise non-decreasing in the past decisions $\{\delta_{j}:j<t\}$ and revealed feedback $\{\theta_{j}:j\in\mathcal{I}_{t}\}$ for the GAIF procedure.
Theorem 1 (Online mFDR and FDR control for GAIF).
Let $\{\alpha_{t}\}_{t\in\mathbb{N}}$ be a sequence of test levels satisfying $\widehat{\mathrm{FDP}}_{\rm GAIF}(t)\leq\alpha$. Then:
  • (a)
    If the null $p$-values are conditionally super-uniformly distributed as in (6), then the procedure guarantees
    $\operatorname{mFDR}(t)\leq\alpha\quad\text{for all }t\in\mathbb{N}.$
  • (b)
    If the null $p$-values are mutually independent and also independent of the non-null $p$-values, and if the test levels $\{\alpha_{t}\}$ are a monotone sequence of functions of the past for all $t$, then the procedure satisfies
    $\operatorname{FDR}(t)\leq\alpha\quad\text{for all }t\in\mathbb{N}.$
The test levels of GAIF in (5) inherit the monotonicity structure from LORD++, and therefore satisfy the monotonicity condition needed for Theorem 1 (b). The proof of Theorem 1 is provided in Appendix F.2.
3 Extensions of GAIF
In this section, we discuss two important extensions of GAIF. We first develop the Adaptive GAIF framework in Section 3.1, which further enhances GAIF by enabling adaptive $\alpha$-wealth allocation. We then extend our methods to the settings with local dependence in Section 3.2.
3.1 Adaptive GAIF: improve GAIF via adaptive $\alpha$-wealth allocation
Section 2.2 shows that feedback improves power by sharpening the FDP estimate. In online testing, power also depends on how the available $\alpha$-wealth is allocated over time, since the test levels $\alpha_{t}$ must be chosen sequentially. This motivates an improvement direction beyond feedback-enhanced FDP estimation, namely adaptive $\alpha$-wealth allocation.
To motivate this idea, we revisit the difference between LORD++ and SAFFRON. The SAFFRON uses a weighting function $\kappa(p):=\frac{\mathbb{I}\{p>\lambda\}}{1-\lambda}$ in its FDP estimator (2). This construction is typically interpreted as a way to estimate the null proportion in the online FDR literature, and has been studied in related contexts, notably in the offline multiple testing setting [Storey et al., 2004, Döhler and Meah, 2023].
Here we emphasize another complementary perspective: $\kappa(\cdot)$ also serves as a guiding mechanism for adaptive $\alpha$-wealth allocation. Indeed, under LORD++, the test levels satisfy $\alpha_{t}\leq\alpha\sum_{j=1}^{t}\delta_{j}-\sum_{j=1}^{t-1}\alpha_{j}$. By contrast, SAFFRON enforces
$\alpha_{t}\leq\alpha\,\frac{1-\lambda}{\mathbb{I}\{p_{t}>\lambda\}}\sum_{j=1}^{t}\delta_{j}-\frac{1}{\mathbb{I}\{p_{t}>\lambda\}}\sum_{j=1}^{t-1}\alpha_{j}\mathbb{I}\{p_{j}>\lambda\}.$ (7)
When $p_{t}\leq\lambda$, the denominator in (7) vanishes, and the upper bound is interpreted as $\infty$, allowing more aggressive investment at promising tests. When $p_{t}>\lambda$ and $p_{j}>\lambda$ for all $j\leq t-1$, the upper bound becomes $\alpha(1-\lambda)\sum_{j=1}^{t}\delta_{j}-\sum_{j=1}^{t-1}\alpha_{j}$, which is smaller than the LORD++ bound. Thus, the weighting function $\kappa(\cdot)$ effectively spends less $\alpha$-wealth on unpromising tests and reserves more for promising ones.
This analysis suggests that the GAIF framework in Section 2.2 can be further improved by adopting an adaptive $\alpha$-wealth allocation strategy. Next, we demonstrate that explicitly incorporating observed $p$-value patterns enables more efficient $\alpha$-wealth allocation. We formalize this idea below:
Definition 3.1 (Adaptive GAIF).
Define the adaptive feedback-enhanced FDP estimator as
@@ 3.1 Adaptive GAIF: improve GAIF via adaptive $\alpha$-wealth allocation
$\widehat{\mathrm{FDP}}_{\mathrm{Ada\text{-}GAIF}}(t):=\frac{\sum_{j\in\mathcal{I}_{t}}(1-\theta_{j})\,\alpha_{j}\kappa(p_{j})\;+\;\sum_{j\in\bar{\mathcal{I}}_{t}}\alpha_{j}\kappa(p_{j})}{1\vee\sum_{j=1}^{t}\delta_{j}},$
where $\kappa(p)=\frac{\mathbb{I}\{p>\lambda\}}{1-\lambda}$ and $\lambda\in[0,1]$ is a user-chosen parameter for identifying large pp‑values. A procedure is called an Adaptive GAIF procedure if it assigns test levels $\{\alpha_{t}\}$ such that, for every $t$,
$\widehat{\mathrm{FDP}}_{\mathrm{Ada\text{-}GAIF}}(t)\leq\alpha.$ (8)
The key difference between GAIF and Adaptive GAIF is the introduction of the $\kappa(p)$ in the FDP estimate. This weight returns the unused $\alpha$-budget from tests with large $p$-values (i.e., $p_{j}>\lambda$) for future allocation, while feedback continues to reduce the slack in FDP estimation through the revealed $\theta_{j}$’s. If feedback is ignored so that $\mathcal{I}_{t}=\emptyset$, the FDP estimator of Adaptive GAIF reduces to that of SAFFRON.
@@
In principle, other constructions of $\kappa(\cdot)$ may further optimize $\alpha$-wealth allocation while ensuring online FDR control. This online optimization remains an open problem and a direction for future work [Gang et al., 2023]. Instead, we focus on the SAFFORN-type weight $\frac{\mathbb{I}\{p>\lambda\}}{1-\lambda}$, which offers a theoretically grounded and practically appealing compromise with robust empirical performance.
We next present concrete instances of Adaptive GAIF. Its thresholds $\{\alpha_{t}\}$ are in the same spirit as SAFFRON, but with the tightened, feedback-aware constraint in (8). For $t=1$, $\alpha_{1}^{\text{Ada-GAIF}}=\min\{\gamma_{1}s_{0},\lambda\}$. For $t>1$,
$\alpha_{t}^{\text{Ada-GAIF}}=\min\bigl\{\lambda,s_{0}\gamma_{t-C_{0+}}+\{(1-\lambda)\alpha-s_{0}\}\gamma_{t-\tau_{1}-C_{1+}}+(1-\lambda)\alpha\sum_{j\geq 2}\gamma_{t-\tau_{j}-C_{j+}}+\sum_{j:j\in\mathcal{I}_{t}}\gamma_{j}\alpha_{j}\theta_{j}\bigr\},$ (9)
where $C_{j+}=C_{j+}(t)=\sum_{i=\tau_{j}+1}^{t-1}C_{i}$ and $C_{t}=\mathbb{I}\{p_{t}\leq\lambda\}$.
The Adaptive GAIF procedures benefit simultaneously from feedback-enhanced FDP estimation and adaptive $\alpha$-wealth allocation, and therefore are expected to achieve higher power. We denote the concrete Adaptive GAIF procedures under full and instant feedback, bandit and instant feedback, full and delayed feedback, and bandit and delayed feedback by SF, SF-BI, SF-FD, and SF-BD, respectively.
@@
Consider LF of GAIF and SF of Adaptive GAIF under the full and instant feedback scenario. Figure 1 depicts the resulting testing thresholds $\{\alpha_{t}\}$ over time $t$ for various procedures applied to Gaussian observations. It is clear that our methods yield larger thresholds after improving the gap via feedback, with $\alpha_{t}^{\text{Ada-GAIF}}>\alpha_{t}^{\text{SAFFRON}}$ and $\alpha_{t}^{\text{GAIF}}>\alpha_{t}^{\text{LORD++}}$ on average. This illustrates that the GAIF and Adaptive GAIF procedures leverage the available $\alpha$-wealth more effectively through the consideration of feedback.
@@
Figure 1: Results for testing Gaussian observations. Line charts depict the average of $\alpha_{t}$ over time $t$ under various testing procedures, based on 500 replications with a non-null proportion of $\pi_{1}=0.5$. Details of the data generation process are provided in Section 5.1.
We now establish the theoretical results for Adaptive GAIF. Similar to (6), the conditional super-uniformity condition on the null $p$-values is
$\Pr\bigl(p_{t}\leq\alpha_{t}\,\bigm|\,\mathcal{J}_{t-1}\bigr)\;\leq\;\alpha_{t}\quad\text{for all }t\in\mathcal{H}_{0},$ (10)
where $\mathcal{J}_{t-1}:=\sigma(\delta_{1:t-1};C_{1:t-1};\{\theta_{j}\}_{j\in{\mathcal{I}}_{t}})$ is the enlarged filtration, where $C_{j}=\mathbb{I}\{p_{j}\leq\lambda\}$. And we extend the monotonicity definition in Theorem 1 by additionally requiring $\alpha_{t}$ to be coordinate-wise non-decreasing in the indicators $\{C_{j}:j<t\}$.
Theorem 2 (Online mFDR and FDR control for Adaptive GAIF).
Let $\{\alpha_{t}\}_{t\in\mathbb{N}}$ be a sequence of test levels satisfying $\widehat{\mathrm{FDP}}_{\rm Ada\text{-}GAIF}(t)\leq\alpha$. Then:
@@
    If the null $p$-values are conditionally super-uniform (10), the procedure guarantees
    $\operatorname{mFDR}(t)\leq\alpha\quad\text{for all }t\in\mathbb{N}.$
@@
    If the null $p$-values are mutually independent and independent of the non-null $p$-values, and if the test levels $\{\alpha_{t}\}$ form a monotone sequence of functions of the past for all $t$, then the procedure satisfies
    $\operatorname{FDR}(t)\leq\alpha\quad\text{for all }t\in\mathbb{N}.$
The test levels of Adaptive GAIF in (9) inherit the monotonicity structure from SAFFRON, and therefore satisfy the monotonicity condition needed for Theorem 2 (b). The proof of Theorem 2 is provided in Appendix F.3.
3.2 GAIF under local dependence
This subsection focuses on extending the GAIF and Adaptive GAIF framework to settings with local dependence (as defined in Definition 3.2). This extension is useful in applications where nearby tests may be dependent, while long-range dependence is absent or weak. For a strategy that addresses more general dependence structures via e-values, the reader may refer to Appendix B.2.